A body of400g slides on a rough horizontal surface. If the frictional force is 3.0, the magnitude of contact force is

Grade:8

1 Answers

Sarfaraz Ahamed
6 years ago
The answer is 5 newton
The contact force on the body =(N+ f)1/2
N=Normal Force on the body which is 0.4kg*10m/s2=4Newton
f=frictional force is 3 Newton (given)
therefore Contact Force is (42 + 32)1/2 = 5 Newton